OpenStudy (anonymous):

This is a general question but how does 11/18 become 11:7. How do you change fraction into odds and vice versa? Does the numerator go to the left and denominator - numerator right?

OpenStudy (anonymous):

you are going from probabilities to odd right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{11}{18}\] is a probability

OpenStudy (anonymous):

gamblers use odd so they interpret \(\frac{11}{18}\) like this: out of every 18 games, i win 11, lose 7 odds in favor of me winning is \(11:7\) and odds against are \(7:11\)

OpenStudy (anonymous):

Ohh okay I was a bit confused thanks for the explanation.

OpenStudy (anonymous):

yw the less intuitive but more mathematical explanation is if the probability i win is \(p\) then the probability i lose is \(1-p\) the odds on favor of me winning is therefore \[p:1-p\] or \[\frac{p}{1-p}\]